# First Step Towards Matter Power Spectrum Reconstruction with Stage-III Weak Gravitational Lensing Surveys

### v1

@@ 3 Data
We use the latest public data releases of the two-point statistics of three Stage-III WL surveys: the Kilo-Degree Survey data release 4 (KiDS-1000; Kuijken et al., 2019), the Dark Energy Survey year 3 (DES year 3; Abbott et al., 2022), and the Hyper Suprime-Cam Year 3 (HSC year 3; Dalal et al., 2023). Collectively, these surveys are referred to as Stage-III WL surveys. We use each survey’s summary statistics, source redshift distributions, and full covariance matrix, as described in Asgari et al. (2021); Hildebrandt et al. (2020); Joachimi et al. (2021) (KiDS-1000); Abbott et al. (2022); Myles et al. (2021); Friedrich et al. (2021) (DES year 3); and Shirasaki et al. (2019); Li et al. (2023); Rau et al. (2023) (HSC year 3). We cut the data vectors at the same angular scales as the fiducial survey analyses. In Table 1, we summarize the main characteristics of the surveys and the statistics we use in this analysis. The table lists the sky area covered, effective source number density ($n_{\mathrm{eff}}$), standard deviation of the galaxy shape measurement error, and number of tomographic bins. As indicated in the final column, for KiDS-1000, we use the 2PCFs, COSEBIs, and band powers, whereas for HSC and DES year 3, we only use the 2PCFs. The former allows us to quantify the difference between the same measurements from a single survey and the scales that are probed by different statistics, whereas using all statistics allows us to quantify the variations between different surveys.
Table 1: Survey characteristics. For each survey, the sky area covered; total effective source number density, $n_{\mathrm{eff}}$; the standard deviation of the error of galaxy shape measurements, $\sigma_{e}$; the number of tomographic bins; and the statistics considered in this work.
@@
KiDS-1000 1006 6.2 0.265 5 $\xi_{\mathrm{+/-}}$, $E_{n}$, $\mathcal{C}_{\mathrm{E},l}$
@@
HSC year 3 416 15 0.236 4 $\xi_{\mathrm{+/-}}$
@@ 4 Methodology
Section 2 has provided the formalism to estimate the cosmic shear observables from the matter power spectrum. Here, we provide the model we use in our analysis.
For the matter power spectrum, we assume a double power law in wavenumber ($k$) and scale factor ($a$) as
@@
where the amplitude $A$, and exponents $p$ and $m$ are the parameters that will be fitted for. The pivot points $k_{\mathrm{piv}}$ and $a_{\mathrm{piv}}$ are free as they rescale the amplitude. We will use two sets of pivots: first, we determine the optimal values of the pivot points for each survey and statistic by finding the values of $k_{\mathrm{piv}}$ and $a_{\mathrm{piv}}$ for which the covariance between the amplitude $A$ and exponents $p$ and $m$ is zero. This method provides the power spectrum amplitude of the point on the ($k,a$) plane most tightly constrained by the statistic. We will provide the best-fit parameters of the double power law parameters assuming these pivot points. After, to more straightforwardly compare the three surveys, we adopt a common pivot for all surveys: $k_{\mathrm{piv}}=0.5\,h/$Mpc and $a_{\mathrm{piv}}=0.75$. We estimate the best-fit parameters that maximize the log-likelihood using the public MCMC sampler emcee (Foreman-Mackey et al., 2013). We use flat priors in the range $\log_{\mathrm{10}}A\in[0,10],\,p\in[-2,0.5]$, and $m\in[-5,5]$. We compute the reduced chi-square value between the double power fits and measurements to quantify the goodness of fit.
@@ 5 Results
5.1 Double power law constraints
We first list the results of determining the optimal pivot points for each survey. Table 2 lists the best-fit pivot points in wavenumber and scale factor for each of the 5 statistics considered in this work and for the choice of $A_{\mathrm{IA}}=0$ or 1 in the third and fourth column, respectively.
Table 2: For each statistic and choice of intrinsic alignment amplitude ($A_{\mathrm{IA}}$) the values of the pivot points that minimize the covariance between the amplitude $\log_{10}A$, and exponents $m$ and $p$ and the best-fit parameters for the double power-law parameters; amplitude, $\log_{10}A$; wavenumber exponent, $p$; scale factor exponent, $m$. The final column indicates the reduced chi-square value of the double power law fit, $\chi_{\mathrm{red}}^{2}$.
Statistic $A_{\mathrm{IA}}$ $k_{\mathrm{piv}}\,[h/$Mpc] $a_{\mathrm{piv}}$ $\log_{10}A\,[\mathrm{Mpc}^{3}/h^{3}]$ $p$ $m$ $\chi_{\mathrm{red}}^{2}$
KiDS-1000 COSEBIs 0 1.20 0.74 $2.04\pm 0.03$ $-1.30\pm 0.04$ $\phantom{-}0.3\pm 0.8$ 1.25
KiDS-1000 COSEBIs 1 1.24 0.75 $2.03\pm 0.03$ $-1.30\pm 0.04$ $\phantom{-}0.2\pm 0.9$ 1.26
KiDS-1000 band powers 0 1.03 0.74 $2.16\pm 0.03$ $-1.25\pm 0.06$ $-0.1\pm 1.0$ 1.39
KiDS-1000 band powers 1 0.93 0.74 $2.21\pm 0.03$ $-1.26\pm 0.06$ $-0.1\pm 1.1$ 1.40
KiDS-1000 2PCFs 0 3.64 0.74 $1.44\pm 0.04$ $-1.24\pm 0.03$ $\phantom{-}0.0\pm 0.8$ 1.25
KiDS-1000 2PCFs 1 4.75 0.74 $1.30\pm 0.04$ $-1.24\pm 0.03$ $-0.3\pm 0.8$ 1.25
DES year 3 2PCFs 0 2.17 0.79 $1.92\pm 0.03$ $-1.14\pm 0.02$ $\phantom{-}1.0\pm 0.5$ 1.14
DES year 3 2PCFs 1 2.01 0.79 $1.97\pm 0.03$ $-1.14\pm 0.02$ $\phantom{-}1.0\pm 0.5$ 1.14
HSC year 3 2PCFs 0 0.80 0.71 $2.36\pm 0.03$ $-1.27\pm 0.03$ $-0.2\pm 0.4$ 1.40
HSC year3 2PCFs 1 0.78 0.71 $2.28\pm 0.03$ $-1.27\pm 0.03$ $-0.3\pm 0.4$ 1.38
@@ 5.1 Double power law constraints
Table 2 also includes the best-fit parameter values and their 16th and 84th percentiles of the double power law fit. For each fit, the parameter best-fit values correspond to an inference assuming the listed pivot point for that statistic and the value of intrinsic alignment amplitude. The best-fit value of the amplitude parameter may, therefore, differ as this is the value of the double power-law model at ($k,a$) = $(k_{\mathrm{piv}},a_{\mathrm{piv}}$). In Appendix A, we show an example of a double fit power law prediction to the KiDS-1000 COSEBIs.
The 3 KiDS-1000 statistics generally find the same values for the power law exponents $m$ and $p$, which are consistent with the HSC year 3 2PCFs estimates. The DES year 3 2PCFs prefer a weaker dependence on wavenumber and a stronger dependence on scale factor.
@@
Stage-III WL surveys primarily probe the $\Lambda$-dominated era, during which linear growth eventually comes to a halt. However, this is a gradual process, and in addition, non-linear growth continues for longer: predictions obtained with class (Blas et al., 2011) indicate an increase in the power spectrum between redshift 0.5 and 0.1 of a factor of 1.5 (1.9) at wavenumber $k=0.1$ (1.0) $h/$Mpc, corresponding to $m$ values of 1.3 (2.1). Therefore, all data sets we analyze here mildly favor a slower growth than predicted by $\Lambda$CDM. This is most apparent in the HSC year 3 2PCFs, which prefer $m=-0.3\pm 0.4$.
Comparing the three different KiDS-1000 estimates for $p$, we find the 2PCFs provide the tightest constraint. The $1\sigma$ uncertainty on $p$ is a factor two smaller than the uncertainty provided by the band powers estimate. This is not unexpected: for the chosen configuration, the 2PCFs probe the largest range of scales, as illustrated in Fig. 1 of Asgari et al. (2021).
Interestingly, the assumed value for $A_{\mathrm{IA}}$ does not markedly affect the results. There are no significant differences between the assumed value for $A_{\mathrm{IA}}$. The results using $A_{\mathrm{IA}}=0$ or 1 are always consistent, and including the intrinsic alignment does not give a better fit, as indicated by the reduced chi-square value in the final column. The values of the reduced chi-square statistics are reasonable for the model’s simplicity. For example, the model does not capture the BAO wiggles and can still fit the data over a broad wavenumber and redshift range. The reduced chi-square values for the KiDS-1000 COSEBIs are 1.25 ($A_{\mathrm{IA}}=0$) and 1.26 ($A_{\mathrm{IA}}=1$). These values indicate the fits are of equal quality as the best-fit $\Lambda$CDM prediction from Asgari et al. (2021), who quote a value of $\chi_{\mathrm{red}}^{2}=1.2$. The same holds for the KiDS-1000 band powers and 2PCFs.
@@
Figure 2: Posterior contours for a double power law model as indicated in the top. The results correspond to setting the pivot points to $k_{\mathrm{piv}}=0.5\leavevmode\nobreak\ h/$Mpc and $a_{\mathrm{piv}}=0.75$. Left: Assuming no contribution from intrinsic alignment. Right: contours from assuming an intrinsic alignment strength set by $A_{\mathrm{IA}}=1$. The panels show the posterior distribution of the inferences using the KiDS-1000 band powers (black), COSEBIs (green), 2PCFs (yellow), HSC year 3 2PCFs (blue), and DES year 3 2PCFs (red). The results are consistent at the $\sim 2\sigma$ level.
@@ 5.2 Comparison to $\Lambda$CDM predictions
Next, in Fig. 3, we compare our power law prediction to the best-fit $\Lambda$CDM predictions from the KiDS-1000 COSEBIs inference assuming $A_{\mathrm{IA}}=1$. The green curve shows the double power prediction. The thin solid curves are all matter power spectrum predictions from the KiDS-1000 COSEBIs $\Lambda$CDM inference chains generated using class (Blas et al., 2011) and HMCode (Mead et al., 2015). These curves are colored by their value of $S_{8}$, as indicated by the color bar. The solid and dashed gray curves show the best-fit and 16th and 84th percentiles of the $\Lambda$CDM prediction from the KiDS-1000 COSEBIs (Asgari et al., 2021), respectively. The double power law fit is consistent with the $\Lambda$CDM prediction over two decades in wavenumber. At the largest wavenumbers ($k>0.5\leavevmode\nobreak\ h/$Mpc), the double power law overestimates the power compared to the $\Lambda$CDM prediction as well as the largest scales ($k<10^{-2}\leavevmode\nobreak\ h/$Mpc, not shown), which are beyond the scales to which cosmic shear is most sensitive. Within this range, the simplistic double power law is consistent with the KiDS-1000 COSEBIs $\Lambda$CDM prediction. The 16th and 84th percentile range typically covers half a dex, illustrating the matter power spectrum is still poorly constrained.
Figure 3: Matter power spectrum at $z=0.35$. The colored curves show the predictions from the KiDS-1000 COSEBIs chains from Asgari et al. (2021), and they are colored by their value of $S_{\mathrm{8}}$ as indicated by the color bar. The solid and dashed gray curves show the best-fit and 16th and 84 percentiles of the KiDS-1000 inference, respectively. The green solid curve shows the best-fit prediction of the $A_{\mathrm{IA}}=1$ double power-law model assumed in this work.
@@ 5.3 Suppression of the matter power spectrum
Finally, we show the ratio of the matter power spectrum to a Planck DMO matter power spectrum as a function of wavenumber at $z=0.35$ in Fig. 4. The dashed green curve shows the best-fit double power law estimate from this work. The solid gray curves and shaded area show the best-fit and 16th and 84th percentiles of the KiDS-1000 COSEBIs $\Lambda$CDM chains, respectively. The dotted red curve shows the prediction assuming the one-parameter parametric model from Amon & Efstathiou (2022) and setting $A_{\mathrm{mod}}=0.69$, which is the value found by Amon & Efstathiou (2022) to reconcile KiDS-1000 with Planck.
@@
Figure 4: Suppression of the matter power spectrum compared to a DMO prediction at $z=0.35$. The dashed green curve corresponds to the best-fit prediction from the double power to the KiDS COSEBIs with $A_{\mathrm{IA}}=1$. The solid gray curve shows the best fit of the KiDS-1000 $\Lambda$CDM inference from Asgari et al. (2021), and the shaded area covers the 16th to 84th percentile region. The dotted red curve uses the parametric model from Amon & Efstathiou (2022) with $A_{\mathrm{mod}}=0.69$.

### v2

@@ 3 Data
We used the latest public data releases of the two-point statistics of three Stage III WL surveys: the Kilo-Degree Survey data release 4 (KiDS-1000; Kuijken et al., 2019), the Dark Energy Survey year 3 (DES year 3; Abbott et al., 2022), and the Hyper Suprime-Cam Year 3 (HSC year 3; Dalal et al., 2023). Collectively, these surveys are referred to as Stage III WL surveys. We use each survey’s summary statistics, source redshift distributions, and full covariance matrix, as described in Asgari et al. (2021), Hildebrandt et al. (2020), and Joachimi et al. (2021) for KiDS-1000; Abbott et al. (2022), Myles et al. (2021), and Friedrich et al. (2021) for DES year 3; and Shirasaki et al. (2019), Li et al. (2023), and Rau et al. (2023) for HSC year 3. We cut the data vectors at the same angular scales as the fiducial survey analyses. In Table 1 we summarize the main characteristics of the surveys and the statistics we used in this analysis. The table lists the sky area covered, effective source number density ($n_{\mathrm{eff}}$), standard deviation of the galaxy shape measurement error, and number of tomographic bins. As indicated in the final column, for KiDS-1000 we used the 2PCFs, COSEBIs, and band powers, whereas for HSC and DES year 3 we only used the real space 2PCFs from Li et al. (2023) and Abbott et al. (2022), respectively. The former allows us to quantify the difference between the same measurements from a single survey and the scales that are probed by different statistics, whereas using all statistics allows us to quantify the variations between different surveys.
Table 1: Survey characteristics.
@@
KiDS-1000 1006 6.2 0.26 5 $\xi_{\mathrm{+/-}}$, $E_{n}$, $\mathcal{C}_{\mathrm{E},l}$
@@
HSC year 3 416 15 0.24 4 $\xi_{\mathrm{+/-}}$
 [^2] Listed for each survey: sky area covered; total effective source number density, $n_{\mathrm{eff}}$; standard deviation of the error of galaxy shape measurements, $\sigma_{e}$; number of tomographic bins; and the statistics considered in this work.
@@ 4 Methodology
Section 2 provides the formalism for estimating the cosmic shear observables from the matter power spectrum. In this section we provide the model we used in our analysis. For the matter power spectrum, we assumed a double power law in wavenumber ($k$) and scale factor ($a$) as
@@
where the amplitude $A$, and exponents $p$ and $m$ are the parameters that will be fitted for. The pivot points $k_{\mathrm{piv}}$ and $a_{\mathrm{piv}}$ are free as they rescale the amplitude. We used two sets of pivots. First, we determined the optimal values of the pivot points for each survey and statistic by finding the values of $k_{\mathrm{piv}}$ and $a_{\mathrm{piv}}$ for which the covariance between the amplitude $A$ and exponents $p$ and $m$ is zero. This method provides the power spectrum amplitude of the point on the ($k,a$) plane most tightly constrained by the statistic. We provided the best-fitting parameters of the double power-law parameters assuming these pivot points. After, to more straightforwardly compare the three surveys, we adopted a common pivot for all surveys: $k_{\mathrm{piv}}=0.5\,h/$Mpc and $a_{\mathrm{piv}}=0.75$. We estimated the best-fitting parameters that maximize the log-likelihood by using the public Markov chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). We used flat priors in the range $\log_{\mathrm{10}}A\in[0,10],\,p\in[-2,0.5]$, and $m\in[-5,5]$. We computed the reduced chi-square value between the double power fits and measurements to quantify the goodness of fit.
@@ 5 Results
5.1 Double power-law constraints
We first list the results of determining the optimal pivot points for each survey. Table 2 lists the best-fitting pivot points in wavenumber and scale factor for each of the five statistics considered in this work and for the choice of $A_{\mathrm{IA}}=0$ or 1 in the third and fourth column, respectively. The fifth column gives the redshift corresponding to the scale factor pivot point ($z_{\mathrm{piv}}=1/a_{\mathrm{piv}}-1$).
Table 2: Pivot points and best-fitting parameters for the double power-law model.
Statistic $A_{\mathrm{IA}}$ $k_{\mathrm{piv}}\,[h/$Mpc] $a_{\mathrm{piv}}$ $z_{\mathrm{piv}}$ $\log_{10}A\,[\mathrm{Mpc}^{3}/h^{3}]$ $p$ $m$ $\chi_{\mathrm{red}}^{2}$
KiDS-1000 COSEBIs 0 1.20 0.75 0.34 $2.04\pm 0.03$ $-1.30\pm 0.04$ $\phantom{-}0.3\pm 0.8$ 1.25
KiDS-1000 COSEBIs 1 1.24 0.75 0.34 $2.03\pm 0.03$ $-1.30\pm 0.04$ $\phantom{-}0.2\pm 0.9$ 1.26
KiDS-1000 band powers 0 1.03 0.74 0.34 $2.16\pm 0.03$ $-1.25\pm 0.06$ $-0.1\pm 1.0$ 1.39
KiDS-1000 band powers 1 0.93 0.74 0.35 $2.21\pm 0.03$ $-1.26\pm 0.06$ $-0.1\pm 1.1$ 1.40
KiDS-1000 2PCFs 0 3.64 0.74 0.35 $1.44\pm 0.04$ $-1.24\pm 0.03$ $\phantom{-}0.0\pm 0.8$ 1.25
KiDS-1000 2PCFs 1 4.75 0.74 0.35 $1.30\pm 0.04$ $-1.24\pm 0.03$ $-0.3\pm 0.8$ 1.25
DES year 3 2PCFs 0 2.17 0.79 0.26 $1.92\pm 0.03$ $-1.14\pm 0.02$ $\phantom{-}1.0\pm 0.5$ 1.14
DES year 3 2PCFs 1 2.01 0.79 0.26 $1.97\pm 0.03$ $-1.14\pm 0.02$ $\phantom{-}1.0\pm 0.5$ 1.14
HSC year 3 2PCFs 0 0.80 0.71 0.41 $2.36\pm 0.03$ $-1.27\pm 0.03$ $-0.2\pm 0.4$ 1.40
HSC year 3 2PCFs 1 0.78 0.71 0.41 $2.28\pm 0.03$ $-1.27\pm 0.03$ $-0.3\pm 0.4$ 1.38
 [^3] Shown for each statistic and choice of intrinsic alignment amplitude ($A_{\mathrm{IA}}$) are the values of the pivot points that minimize the covariance between the amplitude $\log_{10}A$, and exponents $m$ and $p$ in wavenumber ($k_{\mathrm{piv}}$), scale factor ($a_{\mathrm{piv}}$), and corresponding redshift ($z_{\mathrm{piv}}=1/a_{\mathrm{piv}}-1$) and the best-fitting parameters for the double power-law parameters; amplitude, $\log_{10}A$; wavenumber exponent, $p$; scale factor exponent, $m$. The final column indicates the reduced chi-square value of the double power-law fit, $\chi_{\mathrm{red}}^{2}$.
@@ 5.1 Double power-law constraints
Table 2 also includes the best-fitting parameter values and their 16th and 84th percentiles of the double power-law fit. For each fit the parameter best-fitting values correspond to an inference assuming the listed pivot point for that statistic and the value of intrinsic alignment amplitude. The best-fitting value of the amplitude parameter may therefore differ as this is the value of the double power-law model at ($k,a$) = $(k_{\mathrm{piv}},a_{\mathrm{piv}}$). In Appendix A we show an example of a double-fit power-law prediction to the KiDS-1000 COSEBIs.
The three KiDS-1000 statistics generally find the same values for the power-law exponents $m$ and $p$, which are consistent with the HSC year 3 2PCFs estimates. The DES year 3 2PCFs show a weaker dependence on wavenumber and a stronger dependence on scale factor.
@@
Stage III WL surveys primarily probe the $\Lambda$-dominated era, during which linear growth eventually comes to a halt. However, this is a gradual process; in addition, nonlinear growth continues for a longer time: predictions obtained with class (Blas et al., 2011) indicate an increase in the power spectrum between redshift 0.5 and 0.1 of a factor of 1.5 (1.9) at wavenumber $k=0.1$ (1.0) $h/$Mpc, corresponding to $m$ values of 1.3 (2.1). Therefore, all the data sets we analyze here mildly favor a slower growth than predicted by $\Lambda$CDM. This is most apparent in the HSC year 3 2PCFs, which favor $m=-0.3\pm 0.4$.
Comparing the three different KiDS-1000 estimates for $p$, we find that the 2PCFs provide the tightest constraint. The $1\sigma$ uncertainty on $p$ is a factor of two smaller than the uncertainty provided by the band powers estimate. This is not unexpected: for the chosen configuration, the 2PCFs probe the largest range of scales, as illustrated in Fig. 1 of Asgari et al. (2021).
Interestingly, the assumed value for $A_{\mathrm{IA}}$ does not markedly affect the results. There are no significant differences between the assumed values for $A_{\mathrm{IA}}$. The results using $A_{\mathrm{IA}}=0$ or 1 are always consistent, and including the intrinsic alignment does not give a better fit, as indicated by the reduced chi-square value in the final column. The NLA model assumes that density perturbations are characterized by the Poisson equation. This assumption breaks down under nonlinear gravitational collapse, and therefore the model cannot physically capture the IA signal over the entire range of scales considered in this work (Krause et al., 2016; Asgari et al., 2021). However, as the current constraining power is still limited and the values we adopted for the IA strength correspond to those measured by current surveys, we expect that using a different model, such as the tidal alignment and tidal torquing model (Blazek et al., 2019), or including a redshift (e.g., Joachimi et al., 2011) or luminosity (e.g., Chisari et al., 2016) dependence will not change the insensitivity to IA strength.
The values of the reduced chi-square statistics are reasonable for the model’s simplicity. For example, the model does not capture the BAO wiggles and can still fit the data over a broad range of wavenumbers and redshifts. The reduced chi-square values for the KiDS-1000 COSEBIs are 1.25 ($A_{\mathrm{IA}}=0$) and 1.26 ($A_{\mathrm{IA}}=1$). These values indicate the fits are of equal quality to the best-fitting $\Lambda$CDM prediction from Asgari et al. (2021), who quote a value of $\chi_{\mathrm{red}}^{2}=1.2$. The same holds for the KiDS-1000 band powers and 2PCFs.
Figure 2: Posterior contours for a double power-law model as indicated at the top. The results correspond to setting the pivot points to $k_{\mathrm{piv}}=0.5\leavevmode\nobreak\ h/$Mpc and $a_{\mathrm{piv}}=0.75$. Left: Assuming no contribution from intrinsic alignment. Right: Contours from assuming an intrinsic alignment strength set by $A_{\mathrm{IA}}=1$. The panels show the posterior distribution of the inferences using the KiDS-1000 band powers (black), COSEBIs (green), 2PCFs (yellow), HSC year 3 2PCFs (blue), and DES year 3 2PCFs (red). The results are consistent at the $\sim 2\sigma$ level.
@@ 5.2 Comparison to $\Lambda$CDM predictions
Next, in Fig. 3, we compare our power-law prediction to the best-fitting $\Lambda$CDM predictions from the KiDS-1000 COSEBIs inference assuming $A_{\mathrm{IA}}=1$. The green curve shows the double power prediction. The thin solid curves are all matter power spectrum predictions from the KiDS-1000 COSEBIs $\Lambda$CDM inference chains generated using class (Blas et al., 2011) and HMCode (Mead et al., 2015). These curves are color-coded by their value of $S_{8}$, as indicated by the color bar. The solid and dashed gray curves respectively show the best-fitting and the 16th and 84th percentiles of the $\Lambda$CDM prediction from the KiDS-1000 COSEBIs (Asgari et al., 2021). The double power-law fit is consistent with the $\Lambda$CDM prediction over two decades in wavenumber. At the largest wavenumbers ($k>0.5\leavevmode\nobreak\ h/$Mpc), the double power law overestimates the power compared to the $\Lambda$CDM prediction. The model also overpredicts the power on the largest scales ($k<10^{-2}\leavevmode\nobreak\ h/$Mpc, not shown), which are beyond the scales to which cosmic shear is most sensitive. Within this range, the simplistic double power law is consistent with the KiDS-1000 COSEBIs $\Lambda$CDM prediction. The 16th and 84th percentile range typically covers half a dex, illustrating that the matter power spectrum is still poorly constrained.
Figure 3: Matter power spectrum at $z=0.35$. The colored curves show the predictions from the KiDS-1000 COSEBIs chains from Asgari et al. (2021), and they are color-coded by their value of $S_{\mathrm{8}}$ (see the color bar). The solid and dashed gray curves show the best-fitting and 16th and 84th percentiles of the KiDS-1000 inference, respectively. The green solid curve shows the best-fitting prediction of the $A_{\mathrm{IA}}=1$ double power-law model assumed in this work.
@@ 5.3 Suppression of the matter power spectrum
Finally, we show the ratio of the matter power spectrum to a Planck DMO matter power spectrum as a function of wavenumber at $z=0.35$ in Fig. 4. The dashed green curve shows the best-fitting double power-law estimate from this work. The solid gray curves and shaded area respectively show the best-fitting and the 16th and 84th percentiles of the KiDS-1000 COSEBIs $\Lambda$CDM chains. The dotted red curve shows the prediction assuming the one-parameter parametric model from Amon & Efstathiou (2022) and setting $A_{\mathrm{mod}}=0.69$, which is the value found by Amon & Efstathiou (2022) to reconcile KiDS-1000 with Planck.
@@
Figure 4: Suppression of the matter power spectrum compared to a DMO prediction at $z=0.35$. The dashed green curve corresponds to the best-fitting prediction from the double power law model to the KiDS COSEBIs with $A_{\mathrm{IA}}=1$. The solid gray curve shows the best fit of the KiDS-1000 $\Lambda$CDM inference from Asgari et al. (2021), and the shaded area covers the 16th to 84th percentile region. The dotted red curve uses the parametric model from Amon & Efstathiou (2022) with $A_{\mathrm{mod}}=0.69$.